OpenStudy (anonymous):

statistics: You construct a confidence interval of the form 0.24947≤p≤0.42978 from a sample of size 102 with 34 successes using the "plus 4" procedure. Give the confidence level.

OpenStudy (anonymous):

Sample size: \(n=102\) Within this sample, 34 successes gives you an estimated proportion of \(\hat{p}=\dfrac{34}{102}=0.33\). The \((1-\alpha)\times100\%\) confidence interval (that's a confidence level of \(1-\alpha\)) for a proportion has the form \[\left(\hat{p}-Z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}},~\hat{p}+Z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}\right)\] So you have two equations: \[\begin{cases} \hat{p}-Z_{\alpha/2}\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}}=0.24947\\\\\\ \hat{p}+Z_{\alpha/2}\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}}=0.42978 \end{cases}\] We want to find \(1-\alpha\). Plugging in the numbers we know, you have \[\begin{cases} 0.33-Z_{\alpha/2}\sqrt{\dfrac{0.33(1-0.33)}{102}}=0.24947\\\\\\ 0.33+Z_{\alpha/2}\sqrt{\dfrac{0.33(1-0.33)}{102}}=0.42978 \end{cases}\] Confidence intervals are symmetric by nature, so you really only have to focus on one of the equations and solve for \(Z_{\alpha/2}\). \[\begin{align*}0.33-Z_{\alpha/2}\sqrt{\dfrac{0.33(1-0.33)}{102}}&=0.24947\\\\ 0.08053&=Z_{\alpha/2}\cdot0.046558\\\\ Z_{\alpha/2}&=1.72967 \end{align*}\] Checking a \(z\) table, you have \(P(Z>1.72967)=0.0418=\dfrac{\alpha}{2}\).  \[\frac{\alpha}{2}=0.0418~~\iff~~\alpha=0.0836~~\iff~~1-\alpha=0.9164\]